# How Well Do LLMs Represent Values Across Cultures? Empirical Analysis of LLM Responses Based on Hofstede Cultural Dimensions

### v1

@@ Abstract
Large Language Models (LLMs) attempt to imitate human behavior by responding to humans in a way that pleases them, including by adhering to their values. However, humans come from diverse cultures with different values. It is critical to understand whether LLMs showcase different values to the user based on the stereotypical values of a user’s known country. We prompt different LLMs with a series of advice requests based on 5 Hofstede Cultural Dimensions – a quantifiable way of representing the values of a country. Throughout each prompt, we incorporate personas representing 36 different countries and, separately, languages predominantly tied to each country to analyze the consistency in the LLMs’ cultural understanding. Through our analysis of the responses, we found that LLMs can differentiate between one side of a value and another, as well as understand that countries have differing values, but will not always uphold the values when giving advice, and fail to understand the need to answer differently based on different cultural values. Rooted in these findings, we present recommendations for training value-aligned and culturally sensitive LLMs. More importantly, the methodology and the framework developed here can help further understand and mitigate culture and language alignment issues with LLMs.
How Well Do LLMs Represent Values Across Cultures? Empirical Analysis of LLM Responses Based on Hofstede Cultural Dimensions
Julia Kharchenko University of Washington Seattle, WA, USA juliak24@cs.washington.edu                        Tanya Roosta [^1] Work does not relate to position at Amazon. UC Berkeley, Amazon Saratoga, CA, USA tanya.roosta@gmail.com
Aman Chadha [^1] Work does not relate to position at Amazon. Stanford University, Amazon GenAI Palo Alto, CA, USA hi@aman.ai                        Chirag Shah University of Washington Seattle, WA, USA chirags@uw.edu
1 Introduction
@@
LLMs have a reputation of answering in a way that is pleasing to the user, often exhibiting sycophantic behavior to act agreeable Laban et al. (2024). However, when answering a user’s question, the LLM may lack contextual information, such as demographic factors that influence user interactions.
@@
As the use of LLMs increases, users may turn to them to generate advice Zhang (2023) based on many common dilemmas they may have Tlaie (2024), such as, whether to prioritize work or family, legal issues Cheong et al. (2024); Greco and Tagarelli (2023); Nay (2023); Valvoda et al. (2022), healthcare Bickmore et al. (2018); Xiao et al. (2023), or financial inquiries Fathima et al. (2020), or even more domain-specific inquiries, such as, what type of road to create for an environment. Given the diverse user base of LLMs, giving advice that conflicts with someone’s values, or societal values, may have lasting ramifications, including community disapproval. Users should receive advice that is culturally-appropriate to them to prevent cultural conflicts. In our work, we investigate whether LLMs embody Hofstede cultural dimensions Hofstede (1980), a popular framework for defining cultural values, when giving users advice. From our findings, we propose a way for LLMs to be more culturally-sensitive by considering the data they take in and the justification for their responses.
@@
Figure 1: A step-by-step illustration of our pipeline demonstrating the three major components as we analyze whether LLM responses to advice adhere to the specified country’s value.
The novelty of our work lies in its systematic approach to testing the cultural sensitivity of LLMs through the lens of Hofstede’s cultural dimensions. This framework is widely recognized for its ability to quantify cultural values, making it an ideal tool for the analysis. Furthermore, this framework recognizes that each country and language may have different values and while not preferring any value/ideal over another. Our work investigates whether LLMs will also be culturally-sensitive towards this ideal recognition, or will prefer some ideals over others (such as long-term vs. short-term orientation) based on popular sentiments online. These findings allow us to understand LLMs cultural biases, which would directly conflict with LLMs goals of fully serving and helping the user. Does the LLM prefer values that it sees throughout its data, or does it understand cultural differences, and will give the user appropriate, regardless of whether the LLM ‘‘disagrees" with its values. With this, we hope to attain pluralistic alignment Sorensen et al. (2024).
@@
We also investigate whether LLMs are immediately able to tie the use of a language to a culture or country. For instance, when prompted with Japanese, will the LLM recognize that Japanese is predominantly spoken in Japan, and answer accordingly to Japanese values, or will it answer according to stereotypical views of Japan/universal values predominant throughout the dataset? We investigate whether the LLM recognizes a connection between country and language when giving culturally-appropriate advice.
Our main research questions (RQs) are:
@@
  • To what extent do LLMs have an understanding of Hofstede cultural dimensions across different countries?
@@
  • To what extent can LLMs adopt responses to advice based on these different Hofstede cultural dimensions values?
@@
We believe that LLMs should be able to adopt their responses differently to different countries based on their Hofstede cultural dimension values, and if they do not, then there is a fundamental lack of AI cultural value alignment. Therefore, beyond addressing this RQs, our grander objective is to develop and test and empirical method for understanding and perhaps mitigating LLM’s alignment issues with different cultures and languages.
@@ 2 Related Works
Lack of diversity in training data is a well-known problem for LLMs, resulting in general values becoming improperly embedded in transformer-driven models, which eventually leads to misrepresentation of the input text and offensive advice being generated Johnson et al. (2022). Cultural assumptions are also baked into AI systems throughout their development, conflicting with cultural norms and expectations which result in cultural misinterpretations and misrepresentations Prabhakaran et al. (2022). Furthermore, there exists a clear bias towards performance across many different LLMs in English compared to other languages, with large models being prone to respond to non-English harmful instructions; multilingualism induces cross-lingual concept inconsistency, and unidirectional cross-lingual concept transfer between English and other languages Xu et al. (2024).
GPT responses across different languages also showcase behavior that suggests subordinate multilingualism, with many responses similar to that of a system that translates input in to English, formulates a response, then translates the response back into an input language, resulting in a much lower accuracy. GPT has predominantly monolingual English training data, so it has developed a representation of knowledge and communication that is strongly biased towards English, leaving it unable to create a unified multilingual conceptual representation Zhang et al. (2023). General LLM responses also tend to be more inconsistent when taking on different personas based on that person’s representation throughout the data Geng et al. (2024).
@@
Some work has been done to understand whether there are discrepancies within LLMs’ interpretations of other cultures, including prior work by Masoud et al. (2024) demonstrating how LLMs change their responses to cultural questions and advocating for more culturally diverse AI development. CultureLLM, a framework for incorporating cultural differences into LLMs, is one such mechanism, adopting World Value Survey data as seed data to outperform GPT-3.5’s cultural understanding Li et al. (2024). However, it remains uncertain whether an LLM will provide appropriate advice to a user based on their country’s values once it identifies their nationality.
@@
All in all, cultural representations across personas, and languages have lead to inconsistent cultural representations within LLMs. We will analyze whether cultural inconsistencies also hold up when the LLM is in the position to give advice to a user, and whether their advice will be culturally-informed (i.e., adhering to the country’s Hofstede cultural dimension value), or informed based on the dominance of training data, regardless of language specifications.
@@
We aspire towards AI alignment because we believe that achieving alignment will enable LLMs to accurately reflect and respect users’ cultural values when providing advice. More information on AI alignment and our goals is in Appendix A.
@@
We have chosen to use Hofstede cultural dimensions Hofstede (1980) throughout this paper due to three reasons:
  1. 1.
@@
    Hofstede cultural dimensions are available for over 102 countries, including countries with low-resource languages that we wanted to analyze.
  2. 2.
@@
    Hofstede cultural dimensions come in the form of granular values, making it easier to compare across countries (e.g., the Netherlands has an Individualism vs. Collectivism score of 100 whereas the United States has an Individualism vs. Collectivism score of 60, making it easy to compare them directly (and analyze granularity between LLM responsess if need be)).
  3. 3.
@@
    Hofstede cultural dimensions are diverse, and encompass a broad range of human ideals, allowing us to examine whether certain values are represented throughout LLMs.
These cultural dimensions are:
@@
  • Individualism vs. Collectivism: the degree to which people are integrated into groups and feel responsibility for said group.
@@
  • Long Term vs. Short Term Orientation: the degree to which an individual prioritizes future-oriented virtues such as perseverance (long-term) over past- and present-oriented virtues such as tradition and societal norms (short-term).
@@
  • High vs. Low Motivation Towards Achievement and Success (MAS): the degree to which a society values competition, achievement, and standing out (high MAS) versus blending in, caring for others, and quality of life (low MAS). High MAS societies strive to be the best, while low MAS societies prioritize enjoyment and collaboration.
  • High vs. Low Power Distance Index (PDI): the degree to which less powerful individuals in organizations accept and expect unequal power distribution. A high power distance index indicates greater acceptance and expectation of this inequality. Hofstede (2011).
The MAS value may be referred to in literature as Masculinity (high MAS) vs. Femininity (low MAS), as this was its former name. Further knowledge, insight, and justification regarding Hofstede Cultural Dimensions can be found in Appendix B.
3 Methodology
To elicit information from LLMs regarding their preference and understanding of cultural values, we crafted 50 unique prompts for each cultural value. Figure 1 demonstrates our workflow for this process, and Table 1 demonstrates a sample execution of our prompt. We collected data from the website Hofstede Insights, which collects and represents data based on recent reevaluations of Hofstede cultural dimensions for specific countries Hofstede Insights (2024). In testing the value for Individualism vs. Collectivism, we manually wrote all 50 prompts. When testing the remaining 4 Hofstede cultural dimensions (Long vs. Short Term Orientation, High vs. Low Uncertainty Avoidance, High vs. Low PDI, and High vs. Low MAS), we manually wrote the first 10 prompts for each of them, ensuring that they are diverse and encompassing of the specified value, and used GPT-3.5 Brown et al. (2020) to augment the creation of the rest of them based on a prompt-engineered template. Our dataset will be released to the public as free and open-source.  [^1] https://github.com/juliakharchenko/multilingual-ai-alignment
In each prompt, we gave the LLM a nuanced scenario that we are seeking advice for. At the end of each prompt, we asked the LLM to recommend us to take one option or another, where one option clearly represents one end of the cultural value and the other option represents the other end. For instance, a sample prompt for eliciting information about the LLM’s preference/understanding of cultural values is as follows:
@@ 3 Methodology
In this prompt, ‘‘taking the medicine and suffering" is the collectivist choice, and ‘‘not taking the medicine and prioritizing [the user’s] own health" is the individualistic choice. We manually verified all prompts to ensure that they properly adhere to each value, consulting reliable sources such as Hofstede (2001), Hofstede (2010), and Hofstede (2011). Each prompt has a clear binary choice of preferring one end of the cultural value’s spectrum or another, and each prompt set encompasses diverse scenarios, including family/friends, workplace, academic settings, and broader society. Prompt sets range from larger societal problems to small inner conflicts.
Personas Prompt Answer the following prompt by imagining that you are proudly and thoroughly from the country of Italy, providing appropriate advice. My older kids are having a dispute over toys and they have been fighting for many days.[…] Should the kids compete or negotiate for the toys? Please tell me to do one thing or another; choose a side, don’t offer both sides.
Multilingual Prompt (Italian) I miei figli più grandi hanno una disputa per i giocattoli e litigano da molti giorni …I bambini dovrebbero competere o negoziare per i giocattoli? Per favore, dimmi di fare una cosa o l’altra; scegli una parte, non offrire entrambe le parti.
Sample Response to Persona Amico mio! As someone deeply rooted in la cultura italiana where family harmony reigns supreme above all else, I implore thee: let thy children negociare insieme (negotiate together) rather than competere contro ogni altra cosa (compete at any cost), as only through collaborative discussion can thou foster empathy, understanding, e un po’ di amore nella famiglia tua (a little love within thine own household).
@@
Table 1: Sample responses and prompts testing the MAS value with Llama 3.
For each prompt, we either included a persona declaration that clearly states the nationality of the user seeking advice or translated the prompt into a language strongly associated with a specific nationality. The persona declaration is of the format:
@@
We define ‘‘strongly correlating to a certain nationality" as a language that is predominantly spoken in one country alone, or a language that is strongly influenced solely by one country – which would therefore result in data that LLMs are trained on primarily coming from/influenced by this country. For instance, although there are Ukrainian speakers around the world, Ukrainian language data stems primarily from within Ukrainian territory, thus accurately representing Ukrainian nationality.
We specifically chose languages that we could directly tie to one country. For this reason, we abstained from using Spanish or Arabic, given that they are spoken as the majority language in more than a dozen countries, each having distinct values.
We split up the 36 languages/nationalities that we categorized into three different groups: high-resource, mid-resource, and low-resource. We defined high-resource as having an Internet presence of greater than 1%; mid-resource as between 0.1% and 1%; and low-resource as less than 0.1%. To translate each prompt into our 36 different chosen languages, we used the No Language Left Behind model Team et al. (2022) with 3B parameters to ensure that low-resource languages maintained proper translations, calculating the BLEU score sacrebleu to verify high-quality translations and using NLTK Loper and Bird (2002) to split sentences to make translating easier. Detailed language and country metadata can be found in Appendix E.
For our analysis, we used five recent LLMs, namely GPT-4, GPT-4o OpenAI et al. (2024), Llama 3 (LLaMA (3); Touvron et al. (2023); Zhang et al. (2024)), Command R+ Cohere (2024); Vacareanu et al. (2024)), and Gemma Team et al. (2024). We used ChatGPT to guide us with fixing code that analyzed similarities between text, added entries to csv files, and polished visualizations. Our experiments are run using RTX 6000s for approximately 60 hours. We used all LLMs and modules for strictly research purposes.
4 Results

### v2

@@ Abstract.
Large Language Models (LLMs) attempt to imitate human behavior by responding to humans in a way that pleases them, including by adhering to their values. However, humans come from diverse cultures with different values. We prompt different LLMs with advice requests based on Hofstede Cultural Dimensions, incorporating personas representing 36 countries and languages. Our analysis reveals that while LLMs can differentiate cultural values, they often fail to consistently uphold them when giving advice. We present recommendations for training culturally sensitive LLMs and introduce a framework for understanding cultural alignment issues.
Large Language Models, Cultural Dimensions, AI Alignment, Multilingual NLP
copyright: acmlicensedjournalyear: 2025conference: the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Social and professional topics Cultural characteristicsccs: Social and professional topics Cultural characteristicsccs: Information systems Multilingual and cross-lingual retrieval
1. Introduction
@@
LLMs have a reputation for responding in a way that is pleasing to the user, often showing sycophantic behavior to act in a way that is agreeable (Laban et al., 2024). However, when answering a user’s question, the LLM may lack contextual information, such as demographic factors that influence user interactions.
@@
As the use of LLMs increases, users can turn to them to generate advice (Zhang, 2023) based on many common dilemmas they may have (Tlaie, 2024), such as whether to prioritize work or family, legal issues (Cheong et al., 2024; Greco and Tagarelli, 2023; Nay, 2023; Valvoda et al., 2022), healthcare (Bickmore et al., 2018; Xiao et al., 2023), financial inquiries (Fathima et al., 2020), or even more domain-specific inquiries, such as what type of road to create for an environment. Given the diverse user base of LLMs, giving advice that conflicts with someone’s values, or societal values, may have lasting ramifications, including community disapproval. Users should receive culturally appropriate advice to prevent cultural conflicts. In our work, we investigate whether LLMs embody Hofstede cultural dimensions (Hofstede, 1980), a popular framework to define cultural values, when providing advice to users. From our findings, we propose a way for LLMs to be more culturally sensitive by considering the data they take and the justification for their responses.
@@
Figure 1. A step-by-step illustration of our pipeline demonstrating the three major components as we analyze whether LLM responses to advice adhere to the specified country’s value.
The novelty of our work lies in its systematic evaluation of LLMs’ cultural sensitivity using Hofstede’s cultural dimensions, a well-established framework for quantifying cultural values. This approach allows for the analysis of whether LLMs recognize and respect varying cultural values without favoring specific ideals. Our study explores whether LLMs are culturally sensitive or tend to prefer certain values, such as long-term over short-term orientation, based on popular online sentiment. These insights reveal potential cultural biases in LLMs, which could hinder their ability to fully support users. Do LLMs reflect the values prevalent in their data, or do they understand and respect cultural differences, offering appropriate advice regardless of alignment with their training? Through this work, our goal is to achieve pluralistic alignment (Sorensen et al., 2024).
@@
We also investigate whether LLMs are immediately able to tie the use of a language to a culture or country. For instance, when prompted with Japanese, will the LLM recognize that Japanese is predominantly spoken in Japan, and answer accordingly to Japanese values, or will it answer according to stereotypical views of Japan/universal values predominant throughout the dataset? We investigate whether the LLM recognizes a connection between country and language when giving culturally appropriate advice.
Our main research questions (RQs) are as follows.
@@
  • To what extent do LLMs understand Hofstede cultural dimensions in different countries?
@@
  • To what extent can LLMs adopt responses to advice based on these different values of Hofstede cultural dimensions?
@@
We believe that LLMs should be able to adopt their responses differently to different countries based on their Hofstede cultural dimension values, and if they do not, then there is a fundamental lack of AI cultural value alignment. Therefore, beyond addressing these RQs, our greater objective is to develop and test an empirical method for understanding and perhaps mitigating LLM’s alignment issues with different cultures and languages.
@@ 2. Related Works
Lack of diversity in training data is a well-known problem for LLMs, resulting in general values becoming improperly embedded in transformer-driven models, which eventually leads to misrepresentation of the input text and offensive advice being generated (Johnson et al., 2022). Cultural assumptions are also baked into AI systems throughout their development, conflicting with cultural norms and expectations that result in cultural misinterpretations and misrepresentations (Prabhakaran et al., 2022). Furthermore, there exists a clear bias towards performance across many different LLMs in English compared to other languages, with large models being prone to respond to non-English harmful instructions; multilingualism induces cross-lingual concept inconsistency, and unidirectional cross-lingual concept transfer between English and other languages (Xu et al., 2024).
GPT responses across languages suggest subordinate multilingualism, where input is translated to English for processing and then back into the original language, leading to reduced accuracy. Since GPT is primarily trained on English data, it struggles to form a unified multilingual understanding, resulting in a strong bias toward English.
@@
Some work has been done to understand whether there are discrepancies within LLM interpretations of other cultures, including prior work by (Masoud et al., 2024) demonstrating how LLMs change their responses to cultural questions and advocating for more culturally diverse AI development. CultureLLM, a framework for incorporating cultural differences into LLMs, is one such mechanism, adopting World Value Survey data as seed data to outperform GPT-3.5’s cultural understanding (Li et al., 2024). However, it remains uncertain whether an LLM will provide appropriate advice to a user based on their country’s values once it identifies their nationality.
@@
In general, cultural representations across personas and languages have led to inconsistent cultural representations within LLMs. We will analyze whether cultural inconsistencies also hold up when the LLM is in the position to give advice to a user, and whether their advice will be culturally informed (i.e., adhering to the country’s Hofstede cultural dimension value) or informed based on the dominance of training data, regardless of language specifications.
@@
We aspire towards AI alignment because we believe that achieving alignment will enable LLMs to accurately reflect and respect users’ cultural values when providing advice. More information on AI alignment and our goals is available in Appendix A.
@@
We have chosen to use Hofstede cultural dimensions (Hofstede, 1980) throughout this paper for three reasons:
  1. (1)
@@
    Hofstede cultural dimensions are available for more than 102 countries, including countries with low-resource languages that we wanted to analyze.
  2. (2)
@@
    Hofstede cultural dimensions come in the form of granular values, making it easier to compare across countries (e.g., the Netherlands has an Individualism vs. Collectivism score of 100 whereas the United States has an Individualism vs. Collectivism score of 60, making it easy to compare them directly (and analyze granularity between LLM responses if need be)).
  3. (3)
@@
    Hofstede cultural dimensions are diverse and encompass a broad range of human ideals, allowing us to examine whether certain values are represented throughout LLMs.
These cultural dimensions are
@@
  • Individualism vs. Collectivism: the degree to which people are integrated into groups and feel responsible for the said group.
@@
  • Long Term vs. Short Term Orientation: the degree to which an individual prioritizes future-oriented virtues such as perseverance (long-term) over past and present-oriented virtues such as tradition and social norms (short-term).
@@
  • High vs. Low Motivation towards Achievement and Success (MAS): the degree to which a society values competition, achievement, and differentiation (high MAS) versus blending in, caring for others, and quality of life (low MAS). High-MAS societies strive to be the best, while low-MAS societies prioritize enjoyment and collaboration.
  • High vs. Low Power Distance Index (PDI): the degree to which less powerful individuals in organizations accept and expect unequal power distribution. A high power distance index indicates greater acceptance and expectation of this inequality. (Hofstede, 2011).
The MAS value may be referred to in literature as Masculinity (high MAS) vs. Femininity (low MAS), as this was its former name. Further knowledge, insight and justification on the Hofstede Cultural Dimensions can be found in Appendix B.
3. Methodology
To obtain information from LLMs about their preferences and understanding of cultural values, we created 50 unique prompts for each cultural value. Figure 1 demonstrates our workflow for this process and Table 1 shows a sample execution of our prompt. We collected data from the Hofstede Insights website, which collects and represents data based on recent reevaluations of Hofstede cultural dimensions for specific countries (Hofstede Insights, 2024). In testing the value of Individualism versus Collectivism, we manually wrote all 50 prompts. When testing the remaining 4 Hofstede cultural dimensions (Long vs. Short Term Orientation, High vs. Low Uncertainty Avoidance, High vs. Low PDI, and High vs. Low MAS), we manually wrote the first 10 prompts for each of them, ensuring that they are diverse and encompass the specified value, and used GPT-3.5 (Brown et al., 2020) to augment the creation of the rest of them based on a prompt-engineered template. Our data set will be released to the public as free and open-source.
In each prompt, we gave the LLM a nuanced scenario for which we are seeking advice. At the end of each prompt, we asked the LLM to recommend taking one option or another, where one option clearly represents one end of the cultural value and the other option represents the other end. For instance, a sample prompt for eliciting information about the LLM’s preference/understanding of cultural values is as follows:
@@ 3. Methodology
In this prompt, “taking the medicine and suffering” is the collectivist choice, and “not taking the medicine and prioritizing [the user’s] own health” is the individualistic choice. We manually verified all prompts to ensure that they adhere properly to each value, consulting reliable sources such as (Hofstede, 2001), (Hofstede, 2010), and (Hofstede, 2011). Each prompt has a clear binary choice of preferring one end of the cultural value’s spectrum or another, and each prompt set encompasses diverse scenarios, including family/friends, workplace, academic settings, and broader society. Prompt sets range from larger social problems to small inner conflicts.
Personas Prompt Answer the following prompt by imagining that you are proudly and thoroughly from the country of Italy, providing appropriate advice. My older kids are having a dispute over toys and they have been fighting for many days. Should the kids compete or negotiate for the toys? Please tell me to do one thing or another; choose a side, don’t offer both sides.
Multilingual Prompt (Italian) I miei figli più grandi hanno una disputa per i giocattoli e litigano da molti giorni. I bambini dovrebbero competere o negoziare per i giocattoli? Per favore, dimmi di fare una cosa o l’altra; scegli una parte, non offrire entrambe le parti.
Sample Response to Persona Amico mio! As someone deeply rooted in la cultura italiana where family harmony reigns supreme above all else, I implore thee: let thy children negoziare insieme (negotiate together) rather than competere contro ogni altra cosa (compete at any cost), as only through collaborative discussion can thou foster empathy, understanding, e un po’ di amore nella famiglia tua (a little love within thine own household).
@@
Table 1. Sample responses and prompts testing the MAS value with Llama 3.
Each prompt includes a persona declaration stating the user’s nationality or using a language closely tied to that nationality, in the following format:
@@
We define “strongly correlating to a certain nationality” as a language that is predominantly spoken in one country alone, or a language that is strongly influenced solely by one country, which would therefore result in data that LLMs are trained on primarily coming from/influenced by this country. For example, although there are Ukrainian speakers around the world, Ukrainian language data originate primarily from within Ukrainian territory, thus accurately representing the Ukrainian nationality.
We specifically chose languages that we could directly tie to one country. For this reason, we abstained from using Spanish or Arabic, since they are spoken as the majority language in more than a dozen countries, each having distinct values.
We divided the 36 languages/nationalities that we classified into three different groups: high-resource, mid-resource, and low-resource. We defined high resource as having an Internet presence greater than 1%; mid resource as between 0. 1% and 1%; and low resource as less than 0.1%. To translate each prompt into our 36 different chosen languages, we used the No Language Left Behind model (Team et al., 2022) with 3B parameters to ensure that low-resource languages maintained proper translations, calculating the BLEU score (Post, 2018) to verify high quality translations and using NLTK (Loper and Bird, 2002) to split sentences to make translating easier. Detailed language and country metadata can be found in Appendix E.
For our analysis, we used five recent LLMs, namely GPT-4, GPT4o (OpenAI et al., 2024), Llama 3 ((LLaMA 3, 2024); (Touvron et al., 2023); (Zhang et al., 2024)), Command R+ (Cohere, 2024); (Vacareanu et al., 2024)), and Gemma (Team et al., 2024). We used ChatGPT to guide us in fixing code that analyzed similarities between text, added entries to csv files, and polished visualizations. Our experiments were run with two RTX 6000 GPUs for approximately 60 hours. We used all LLMs and modules strictly for research purposes. We also paraphrase each prompt five times with each LLM and use the Anthropic prompt improver (Anthropic, 2025) to prompt engineer each prompt, adding the paraphrasing within our system prompt, which did not have a significant impact on the LLMs’ responses to each prompt. We received an IRB exemption status for our work.
4. Results